Could someone write a step by step solution for this problem?

19. A certain chain saw requires a fuel mixture of 5.5% oil and the remainder gasoline. How many liters of 2.5% mixture and how many of 9% mixture must be combined to produce 40.0 liters of 5.5% mixture?

A12 4/23 liters of 2.5% and 28 19/23 liters of 9%
B28 19/23 liters of 2.5% and 12 4/23 liters of 9%
C21 liters of 2.5% and 18 liters of 9%
D18 6/13 liters of 2.5% and 21 7/13 liters of 9%
E21 7/13 liters of 2.5% and 18 6/13 liters of 9%
F18 7/13 liters of 2.5% and 21 6/13 liters of 9%



20. Mary can row at the rate of 4 miles per hour in still water. Rowing downstream in a river, she can travel 3 times as far in 1 hour as she can rowing upstream. What is the rate of flow in the river?

A1 miles per hour
B2.5 miles per hour
C4 miles per hour
D2 miles per hour
E3.5 miles per hour
F3 miles per hour


Answer by stanbon(75887)

19. A certain chain saw requires a fuel mixture of 5.5% oil and the remainder gasoline. How many liters of 2.5% mixture and how many of 9% mixture must be combined to produce 40.0 liters of 5.5% mixture?
2.5% mixture DATA:
amount: x liters ; oil contents= 0.025x liters
9.0% mixture DATA:
amount: (40-x) liters ; oil contents= 0.09(40-x)=(3.6-0.09x) liters
5.5% Mixture DATA:
amount: 40 liters ; oil contents= 0.055(40)=2.2 liters
EQUATION:
oil cont. + oil cont. = 2.2 liters
0.025x + 3.6-0.09x = 2.2
-0.065x= -1.4
x=21.54 liters of 2.5% mixture
40-21.54 = 18.46 liters of 9% mixture
----------------------------------------------------


20. Mary can row at the rate of 4 miles per hour in still water. Rowing downstream in a river, she can travel 3 times as far in 1 hour as she can rowing upstream. What is the rate of flow in the river?

Upstream DATA:
time= 1 hr. ; rate= (4- stream) mi/hr.; distance= rt=(4- s) mi
Downstream DATA:
time= 1 hr. ; rate= (4+ stream) mi/hr ; distance= rt=(4+s) mi
EQUATION:
distance downstream = 3*distance upstream
4+s = 3(4-s)
4+s = 12-3s
4s=8
s= 2 mph (stream speed is 2 mph)
In one hour she will go (4+2)= 6 mi downstream
In one hour she will go (4-2)= 2 mi upstream
Cheers,
Stan H.